### v8

@@ 4.3 Evaluation Criteria
OM typically measures performance using precision, recall, and the F1 score. Given a gold standard reference (R) and a system-discovered alignment (A), Precision measures matching correctness and recall measures matching completeness, while F1 score offers a harmonic mean to balance correctness and completeness.
@@
4.4 Experiment Results
Figure 4 shows the result of evaluating Agent-OM and Agent-OV on the OV testbed. We choose to use the Meta open-source model llama-3-8b for our experiments in this paper, eschewing the higher-end models which offer small improvements that may not justify the financial investment. We use embeddings from the same LLM model to further avoid expensive API calls and thereby provide an entirely free-to-use version of Agent-OV. The hyperparameter settings are similarity_threshold=0.90similarity\_threshold=0.90 and $top@k=3$ across all the system alignments generated. We set $top@k=3$ based on empirical experience that 3-5 works well for Agent-OM. We experiment with the alternative values of the similarity threshold later in this paper.
@@ 4.4 Experiment Results
The results indicate that it is possible to unify the OM and OV tasks using the OM4OV pipeline. Both Agent-OM and Agent-OV can produce alignments with precision, recall, and F1 scores. However, we can see that Agent-OM consistently performs unusually well across all alignments. This is because the traditional OM system captures only matched entities between two ontologies, which in the OV task correspond to the remain and update entities. While remain entities commonly comprise the major proportion of entities over two different versions, the measure of remain entities dominates the overall evaluation and other categories are ignored. We believe this is a common situation that leads to misleading performance measurement when using the OM system in OV tasks. Agent-OV overcomes the skewed measure in Agent-OM. When decomposing the measure into four sub-measures, we can observe more informative matching performance across different OV tasks. In general, we observe that the matching performance is highest in remain, followed by add and delete, and relatively low in update. This trend is consistent across different tracks and ontologies. The measures for remain are typically very close to 100%. The measures for add and delete are generally good. Our system uses LLMs as the backend, and LLMs generally have strong background knowledge to detect non-matched entities. The measures for update show scope for improvement. This is because finding non-trivial alignments and appropriate similarity thresholds is not easy. Additionally, we observe a longer computation time in the Anatomy Track. Although Agent-OV has an optimisation module for the matching candidate selection process (inherited from Agent-OM), it is still insufficient for some OV tasks. There is a need to optimise the pipeline for OV in large-scale ontologies.
Figure 5 studies the effect of varying similarity thresholds in Agent-OV. We change the similarity threshold from 1.00 to 0.50 to evaluate its effect on matching performance. While OV entities are classified into four different categories, each category requires a corresponding sub-measurement. Similar to OM tasks, changes in hyperparameter settings in OV tasks lead to a trade-off between precision and recall. Moreover, the hyperparameter settings also influence the sub-measures. We find that the similarity threshold has no effect on detecting remain entities, but significantly affects the detection of update entities and slightly affects the detection of add and delete entities. Furthermore, detecting update entities and detecting add and delete entities are negatively correlated. Lower similarity thresholds can result in more update entities being detected, while higher similarity thresholds may find more add and delete entities.
@@
We further examine the false mappings produced by Agent-OV. Unlike OM tasks, we find that some of these false mappings are not actually untrue. Several examples are listed below.
{bracketenumerate}
False mappings in OV can result from flawed ontology design choices. In the following example, the reference shows that cmt:writtenBy is updated to cmt:isWrittenBy, but the OV system will consider this entity to remain unchanged because a false mapping (cmt:writtenBy, cmt:writtenBy, $\equiv$, 1) is detected in the remain subset. This issue is caused by inconsistent naming convention practice. In the CMT ontology, cmt:hasAuthor and cmt:writtenBy are different entities because one is targeting the paper (cmt:Paper, cmt:hasAuthor, cmt:Author) and one is targeting the review (cmt:review, cmt:writtenBy, cmt:reviewer). However, using cmt:writtenBy for the paper still makes sense (cmt:Paper, cmt:writtenBy, cmt:Author). Within one ontology, the meaning of two entities is too close to be distinguished and therefore leads to them being synonyms for each other. Ideally, we should avoid this type of ontology design. This example also demonstrates a unique benefit of using OM for OV, which could potentially assist in ontology design.
cmt:hasAuthor => cmt:writtenBy
cmt:writtenBy => cmt:isWrittenBy
@@
cmt:Chairman => cmt:Chair
cmt:ConferenceChair => cmt:General_Chair
False mappings in OV can arise from inappropriate settings of the similarity threshold. In the following example, if similarity_threshold(s)=0.95similarity\_threshold\ (s)=0.95, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If similarity_threshold(s)=0.90similarity\_threshold\ (s)=0.90, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$. Both results are valid because defining the boundary between matching and non-matching is context- and application-dependent. For example, the similarity threshold could be relatively higher in the biomedical domain to ensure that each term is unique, whereas the similarity threshold in the conference domain could be relatively lower to improve the interoperability of terminologies used in research conferences.
s = 0.95, cmt:SubjectArea => None
s = 0.90, cmt:SubjectArea => cmt:Topic
@@ 4.5 Discussion
The experimental results have shown that OM systems can be reused for OV tasks, but they hvae limitations and require necessary extensions.
{bracketenumerate}
Skewed Measurement: OM systems has skewed measurement of OV task because unchanged remain entities dominate in practice. OM measures are reusable, but these measures need to be extended into four sub-measures for add, delete, remain, and update performance. Within each sub-measure, the inherent precision-recall trade-off still holds. Across different sub-measures, they are not independent. This is because each change in a part of an alignment will cause other parts to change accordingly. For example, if a new mapping is found in $A\odot$ or $A\otimes$, then the number of mappings in $A\oplus$ and $A\ominus$ will be reduced by one each, and vice versa.
@@ 5 OM4OV Optimisation
Having established the current pitfalls of OM4OV, we explore optimisations for the pipeline. Often, ontology creators provide cross-references to other ontologies to enhance interoperability and facilitate integration. For example, a cross-reference between the CMT ontology and the Conference ontology is provided alongside the CMT ontology. Reusing the cross-references developed for OM tasks, we propose a novel mechanism to reduce the number of matching candidates and also to improve overall OV performance.
@@
Figure 6: Cross-reference mechanism.
@@
Given a reference ontology ($O_{r}$) with an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), two cross-references between $O$ and $O_{r}$ ($R_{or}$) and between $O^{\prime}$ and $O_{r}$ ($R_{o^{\prime}r}$) are defined as:
$R_{or}$ $=\{(e1,e3,c)|e1,e3\in O_{r}\cap O,s\leq c\leq 1\}$ (5)
$R_{o^{\prime}r}$ $=\{(e2,e3,c)|e2,e3\in O_{r}\cap O^{\prime},s\leq c\leq 1\}$
We can use $R_{or}$ and $R_{o^{\prime}r}$ to infer some known mappings between $O$ and $O^{\prime}$ before performing OV. We call these mappings a prior alignment ($A_{\pi}$). After subsequently performing OV, we have our posterior alignment ($A_{\pi^{*}}$). In this setting, $A_{match}$ in OV can be decomposed into two parts: $A_{match}=A_{\pi}\cup A_{\pi^{*}}$.
@@
$A_{\pi}$ can be directly inferred from the two cross-references $R_{or}$ and $R_{o^{\prime}r}$. The equivalence relation is transitive, so for $e1\in O$, $e2\in O^{\prime}$, and $e3\in O_{r}$, if $e1=e3$ and $e2=e3$ then $e1=e2$. $A_{\pi^{*}}$ aims to detect missing mappings from the cross-reference. None of these mappings would come from any pairwise intersection of $O$, $O_{r}$, and $O^{\prime}$ because $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$ are pre-defined as non-matched entities, and the matched entities in $O\cap O_{r}\cap O^{\prime}$ have already been captured in the $A(\pi)$. As a result, $A_{\pi^{*}}$ can be defined within a smaller scope:
$A_{\pi}$ $=R_{or}\cap R_{o^{\prime}r}=\{(e1,e2,c)|e1,e2\in O\cap O_{r}\cap O^{\prime},s\leq c\leq 1\}$ (6)
$A_{\pi^{*}}$ $=\{(e1,e2,c)|e1\in O\setminus O_{r},e2\in O^{\prime}\setminus O_{r},s\leq c\leq 1\}$
@@
An ontology can have multiple cross-references available. In such cases, the prior reference becomes the union of all known cross-references ($R_{or1}$$R_{orn}$), and the ontology used in the posterior alignment ($O_{ra}$) becomes the union of all reference ontologies ($O_{r1}$$O_{rn}$). Therefore, $A_{\pi}$ and $A_{\pi^{*}}$ in multiple cross-references can be formulated as:
@@
$A_{\pi^{*}}$ $=\{(e1,e2,c)|e1\in O\setminus O_{ra},e2\in O^{\prime}\setminus O_{ra},c\geq s\}\ \text{where}\ O_{ra}=O_{r1}\cup O_{r2}\cup.O_{rn}$
@@
Figure 8 compares the effects of different similarity thresholds in the CMT Ontology versioning. The results of the experiment show that the CR mechanism is less sensitive to the similarity threshold than the original OM4OV pipeline, suggesting that it can be helpful in scenarios where the optimal similarity threshold is unclear or difficult to determine.
Figure 8: Evaluation of the cross-reference (CR) mechanism on different similarity thresholds. We apply a 1-D Gaussian filter for smoothing. The standard deviation for a Gaussian kernel is 1 (i.e. $sigma=1$).
@@ 6 Conclusion
In this paper, we systematically analyse similarities and differences between OM and OV tasks and provide formal definitions and task formulations for OM4OV. The experimental results obtained from Agent-OM and Agent-OV demonstrate that the OM4OV pipeline is workable, though it has several limitations and requires extensions. To tackle these issues, we propose a novel optimisation method that uses a cross-reference mechanism to leverage reference alignments from OM for OV. This method (1) overcomes several pitfalls in using OM for OV tasks, (2) significantly reduces the number of matching candidates, and (3) improves overall OV performance.
While OM has a universal measure, our proposed measures for OV have four sub-measures. We plan to investigate a merged formula for OV sub-measures in the future. For example, using a harmonic mean to combine sub-measures. We also plan to apply the OM4OV pipeline in emerging domain ontologies that require consistent changes over time. For example, the smart building ontologies Brick Schema [4], RealEstateCore [16], and ASHRAE Standard 223P [25].
@@
During the preparation of this work, no GenAI was used to produce the manuscript. However, GenAI was intrinsic to the research reported here and is properly identified throughout the work.

### v9

@@ 4.3 Evaluation Criteria
OM typically measures performance using precision, recall, and the F1 score. Given a gold standard reference (R) and a system-discovered alignment (A), precision measures matching correctness and recall measures matching completeness, while F1 score offers a harmonic mean to balance correctness and completeness. OV can reuse these measures, but they need to be extended into four sub-measures for add, delete, remain, and update performance. While an OV method will trade off performance over each of these sets, an OV application problem may well prioritise accuracy on some over others.
@@
4.4 Results
Figure 4 shows the performance of Agent-OM and Agent-OV on the OV testbed and sets a benchmark for further research progress. The fixed seed for the OV testbed is set to 42. We choose to use the Meta open-source model llama-3-8b for our experiments in this paper, eschewing the higher-end models which offer small improvements that may not justify the financial investment. We use embeddings from the same LLM model to further avoid expensive API calls and thereby provide an entirely free-to-use version of Agent-OV. The hyperparameter settings are similarity_threshold=0.90similarity\_threshold=0.90 and $top@k=3$ across all the system alignments generated. We set $top@k=3$ based on empirical experience that 3-5 works well for Agent-OM. We experiment with the alternative values of the similarity threshold later in this paper.
@@ 4.4 Results
The results indicate that it is possible to unify the OM and OV tasks using the OM4OV pipeline. Both Agent-OM and Agent-OV can produce alignments with precision, recall, and F1 scores. However, we can see that Agent-OM consistently performs surprisingly well across all alignments. This is because the traditional OM system captures only matched entities between two ontologies, which in the OV task correspond to the remain and update entities. Because remain entities typically comprise the major proportion of entities over two different versions, the correctness of remain entities dominates the overall evaluation and other categories are de-emphasised. We believe this is a common situation that leads to misleading performance measurement when using an OM system in OV tasks.
Agent-OV overcomes the skewed measure in Agent-OM. By decomposing the measure into four sub-measures, we can observe more informative matching performance across different OV tasks. In general, we observe that the matching performance is highest in remain, followed by add and delete, and relatively low in update. This trend is consistent across different tracks and ontologies. The measures for remain are typically very close to 100%. The measures for add and delete are generally good. Our system uses LLMs as the backend, and LLMs generally have strong background knowledge to detect non-matched entities. The measures for update show scope for improvement. This is because finding non-trivial alignments and appropriate similarity thresholds is not easy. Additionally, we observe a longer computation time in the Anatomy Track. Although Agent-OV has an optimisation module for the matching candidate selection process (inherited from Agent-OM), it is still insufficient for some OV tasks. There is a need to optimise the pipeline for OV in large-scale ontologies.
Figure 5 studies the effect of varying similarity thresholds for the Mouse ontology in Agent-OV. We change the similarity threshold from 1.00 to 0.80 to evaluate its effect on matching performance. While OV entities are classified into four different categories, each category requires a corresponding sub-measurement. Similar to OM tasks, changes in hyperparameter settings in OV tasks lead to a trade-off between precision and recall. Moreover, the hyperparameter settings also influence the sub-measures. We find that the similarity threshold has no effect on detecting remain entities, but significantly affects the detection of update entities and slightly affects the detection of add and delete entities. Detecting update entities and detecting add and delete entities are negatively correlated, as is expected because a near-miss update alignment will fall back to a pair of non-match entities, one from each ontology. Therefore, lower similarity thresholds can result in more update entities being detected, while higher similarity thresholds may find more add and delete entities.
@@
We further examine the false mappings produced by Agent-OV. Unlike OM tasks, we find that some of these identified false mappings are not actually false. Several examples are listed below. {bracketenumerate}
False mappings in OV can result from flawed ontology design choices. In the following example, the reference shows that cmt:writtenBy is updated to cmt:isWrittenBy, but the OV system will consider this entity to remain unchanged because another mapping (cmt:hasAuthor, cmt:writtenBy, c) leads to a false mapping (cmt:writtenBy, cmt:writtenBy, 1) that is detected in the remain subset. In the CMT ontology, cmt:hasAuthor and cmt:writtenBy are different entities because one is targeting the paper (cmt:Paper, cmt:hasAuthor, cmt:Author) and one is targeting the review (cmt:review, cmt:writtenBy, cmt:reviewer). However, using cmt:writtenBy for the paper still makes sense (cmt:Paper, cmt:writtenBy, cmt:Author). Within one ontology, the meaning of two entities is too close to be distinguished and therefore leads to them being synonyms for each other. Ideally, we should avoid this type of ontology design. This example also demonstrates a unique benefit of using OM for OV, which could potentially assist in ontology design.
(cmt:writtenBy, cmt:isWrittenBy, c)
(cmt:hasAuthor, cmt:writtenBy, c)
@@
(cmt:ConferenceChair, cmt:General_Chair, c)
(cmt:Chairman, cmt:Chair, c)
False mappings in OV can arise from inappropriate settings of the similarity threshold. In the following example, if similarity_threshold=0.95similarity\_threshold=0.95, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If similarity_threshold=0.90similarity\_threshold=0.90, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$. Both results are valid because defining the boundary between matching and non-matching is context- and application-dependent. For example, the similarity threshold could be relatively higher in the biomedical domain to ensure that each term is unique, whereas the similarity threshold in the conference domain could be relatively lower to improve the interoperability of terminologies used in research conferences.
(cmt:SubjectArea, None, 0.95)
(cmt:SubjectArea, cmt:Topic, 0.90)
@@ 4.5 Discussion
The experimental results have shown that OM systems can be reused for OV tasks, but they have limitations and require necessary extensions. The following key points need to be taken into consideration when reusing OM systems for OV tasks. {bracketenumerate}
Skewed Measurement: OM systems exhibit skewed measurement of OV task because unchanged remain entities dominate in practice. OM measures are reusable, but these measures need to be extended into four sub-measures for add, delete, remain, and update performance. Within each sub-measure, the inherent precision-recall trade-off still holds. Across different sub-measures, they are not independent. This is because each change in a part of an alignment will cause other parts to change accordingly. For example, if a new mapping is found in $A\odot$ or $A\otimes$, then the number of mappings in $A\oplus$ and $A\ominus$ will be reduced by one each, and vice versa.
@@ 5 OM4OV Optimisation
Having established the current pitfalls of OM4OV, we explore optimisations for the pipeline. Often, ontology creators provide cross-references to alternative ontologies to enhance interoperability and facilitate integration. For example, a cross-reference between the Mouse ontology and the Human ontology is provided in the OAEI anatomy track. Reusing the cross-references developed for OM tasks, we propose a novel mechanism to reduce the number of matching candidates and also to improve overall OV performance.
@@
Figure 6: Illustration of the cross-reference mechanism.
@@
Given a reference ontology ($O_{r}$), an old version of an ontology ($O$) and a new version of the same ontology ($O^{\prime}$), two cross-references between $O$ and $O_{r}$ ($R_{or}$) and between $O^{\prime}$ and $O_{r}$ ($R_{o^{\prime}r}$) are defined respectively as:
$R_{or}$ $=\{(e_{1},e_{3},c)\mid e_{1},e_{3}\in O_{r}\cap O,s\leq c\leq 1\mbox{ and there is no }(e_{1},e_{3},c^{\prime})\mbox{ with }c^{\prime}>c\}$ (5)
$R_{o^{\prime}r}$ $=\{(e_{2},e_{3},c)\mid e_{2},e_{3}\in O_{r}\cap O^{\prime},s\leq c\leq 1\mbox{ and there is no }(e_{2},e_{3},c^{\prime})\mbox{ with }c^{\prime}>c\}$
We can use $R_{or}$ and $R_{o^{\prime}r}$ to infer some known mappings between $O$ and $O^{\prime}$ before performing OV. We call these mappings a prior alignment ($A_{\pi}$). After subsequently performing OV, we have our posterior alignment ($A_{\pi^{*}}$). In this setting, $A_{match}$ in OV can be decomposed into two parts: $A_{match}=A_{\pi}\cup A_{\pi^{*}}$. $A_{\pi}$ can be directly inferred from the two cross-references $R_{or}$ and $R_{o^{\prime}r}$. By assuming the equivalence relation as extracted from the two cross-references is transitive, we have that $e_{1}\in O$, $e_{2}\in O^{\prime}$, and $e_{3}\in O_{r}$, if $e_{1}=e_{3}$ and $e_{2}=e_{3}$ then $e_{1}=e_{2}$. $A_{\pi^{*}}$ aims to detect missing mappings from the cross-reference. None of these mappings would come from any pairwise intersection of $O$, $O_{r}$, and $O^{\prime}$ because $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$ are pre-defined as non-matched entities, and the matched entities in $O\cap O_{r}\cap O^{\prime}$ have already been captured in the $A(\pi)$. As a result, $A_{\pi^{*}}$ can be defined within a smaller scope.
@@
Prior alignment ($A_{\pi}$) and posterior alignment ($A_{\pi^{*}}$) are defined as:
$A_{\pi}$ $=R_{or}\cap R_{o^{\prime}r}=\{(e_{1},e_{2},c)\mid e_{1},e_{2}\in O\cap O_{r}\cap O^{\prime},s\leq c\leq 1\}$ (6)
$A_{\pi^{*}}$ $=\{(e_{1},e_{2},c)\mid e_{1}\in O\setminus O_{r},e_{2}\in O^{\prime}\setminus O_{r},s\leq c\leq 1\}$
An ontology can have multiple cross-references available. In such cases, the prior reference becomes the union of all known cross-references ($R_{or1}$$R_{orn}$), and the ontology used in the posterior alignment ($O_{ra}$) becomes the union of all reference ontologies ($O_{r1}$$O_{rn}$).
@@
$A_{\pi}$ and $A_{\pi^{*}}$ in multiple cross-references can be formulated as:
@@
$A_{\pi^{*}}$ $=\{(e_{1},e_{2},c)\mid e_{1}\in O\setminus O_{ra},e_{2}\in O^{\prime}\setminus O_{ra},c\geq s\}\ \text{where}\ O_{ra}=O_{r1}\cup O_{r2}\cup.O_{rn}$
@@
Figure 8 compares the effects of different similarity thresholds for the Mouse ontology versioning. The results of the experiment show that the CR mechanism is less sensitive to the similarity threshold than the original OM4OV pipeline in Figure 5, suggesting that it can be helpful in scenarios where the optimal similarity threshold is unclear or difficult to determine.
Figure 8: Evaluation of the cross-reference (CR) mechanism on different similarity thresholds. The solid lines contribute to OV performance without the CR mechanism, while the dotted lines contribute to OV performance with the CR mechanism. We apply a 1-D Gaussian filter for smoothing. The standard deviation for a Gaussian kernel is 1 (i.e. $sigma=1$).
@@ 6 Conclusion
In this paper, we systematically analyse similarities and differences between OM and OV tasks and provide formal definitions and task formulations for OM4OV. The experimental results obtained from Agent-OM and Agent-OV demonstrate that the OM4OV pipeline is workable, though it presents several limitations. To tackle these issues, we propose a novel optimisation method that uses a cross-reference mechanism to leverage reference alignments from OM for OV. This method (1) overcomes several pitfalls in using OM for OV tasks, (2) significantly reduces the number of matching candidates, and (3) improves overall OV performance.
We observe three opportunities for future work. (1) While OM has a universal evaluation measure, our proposed measures for OV have four sub-measures. We plan to investigate a merged formula for OV sub-measures in the future, for example, using a harmonic mean to combine sub-measures. (2) In this paper, we limit the semantic analysis to noting changes in 1-hop subgraph relations that can be either ontology keywords (such as domain, range, and subclass) or domain concepts as object properties. A more complete analysis of structural evolution could employ description logic (DL)-based reasoning for subsumption checking, consistency validation, and detection of semantically equivalent expressions with only syntactic differences. This kind of analysis has not been previously explored in OM and is not addressed in this paper, but could be worthwhile further work. (3) We plan to apply the OM4OV pipeline in emerging domain ontologies that demonstrate regular changes over time. For example, it could be used to track version changes and improve interoperability in smart building ontologies, such as Brick Schema [4] and RealEstateCore [16] for which schema-interoperability is imperative [31].
@@
During the preparation of this work, no GenAI was used to produce the manuscript. However, GenAI is intrinsic to the research reported here and is properly identified throughout the work.